### v1

# Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework

Abstract
A modern chiral potential incorporating the three-body force is adopted to investigate bulk properties, spectra, and nuclear responses of closed-(sub)shell nuclei throughout the nuclear chart within a particle-hole (p-h) renormalized random-phase approximation (RRPA) scheme using a Hartree-Fock (HF) single-particle basis. Our analysis shows that all instabilities induced by the quasiboson approximation (QBA) underlying RPA are removed and an overall better consistency with the experiments is achieved for all observables of the investigated nuclei. The residual discrepancies point out the need of going beyond the p-h space.
Introduction. RPA is one of the most widely adopted methods devoted to the study of nuclear spectroscopy. It can be derived from the equation of motion method [1] or from the linear response theory [2].
It was implemented within many different frameworks, some using non-relativistic (e.g. Refs. [3, 4] ) or relativistic (e.g. Refs. [5, 6]) density functionals, other exploiting the Green’s function formalism (e.g. Refs. [7, 8]). The latter approach allows one to go beyond the harmonic approximation by coupling the RPA modes to complex configurations. Other extensions using Skyrme forces were proposed [9]. The investigations based on modern realistic potentials are few and rather recent [10, 11, 12, 13, 14, 15].
RPA is known to be an extension of Tamm-Dancoff approximation (TDA). In both approaches, the eigenvalue problem is formulated within a p-h or two quasiparticle (2qp) spaces. In TDA, the ground state is the HF vacuum. In RPA, it is assumed to be correlated in the formal derivation of the equations of motion, but is replaced by an uncorrelated HF wavefunction in the actual calculation of the RPA matrix elements. This simplification, known as quasi-boson approximation (QBA), induces instabilities at low excitation energies.
Most of the recipes for obviating this shortcoming were enumerated and developed long ago by Rowe [16, 17]. They consist in reintroducing the correlations into the ground state without spoiling the simple structure of the RPA eigenvalue equations. Few investigations were conducted along this line since then [18, 19, 20]. An important extension was achieved and applied to metal clusters in Refs. [21, 22].
It is worth mentioning a more recent approach using the realistic Argonne v18 potential [23], as well as a dressed (D) RPA study performed within a self-consistent Green’s function scheme [14]. A variety of works aiming to go beyond RPA is presented and discussed in a recent review [24].
Other approaches for computing ground-state correlations are available. We mention an extension of TDA [25], a time-dependent density matrix (TDDM) formalism [26], a many-body perturbative approach [27], in-medium similarity renormalization group (IMSRG) [28] and an equation of motion multiphonon method (EMPM) [29].
The need for restoring the ground-state correlations is also dictated by the fact that HF accounts only for a fraction of the binding energy of all closed-(sub)shell nuclei throughout the periodic table if modern realistic potentials are adopted. This emerges blatantly from the results presented in Ref. [29].
In the present work, we perform a systematic study of bulk properties, nuclear responses, and spectra of a large number of nuclei ranging from 4He to 208Pb. We use the chiral potential N2LOGO(394) [30], incorporating the three-body force, to generate a self-consistent HF basis and then solve the eigenvalue equations within TDA, RPA, and RRPA.
To our knowledge, this is the first ab initio extension of RPA which removes the instabilities induced by the QBA. It accounts for the nuclear bulk properties and provides a unified scheme where energy levels and responses are referred to the true correlated (rather than unperturbed) ground state.
The RRPA formalism. Following the procedure of Refs. [17, 21, 22] we consider states of the form
Here denotes a renormalized p-h creation operator with respect to the HF vacuum, and
where are the particle () and hole () ground-state occupation numbers. The ground state is obtained by imposing the condition
In the QBA underlying RPA, is replaced by the unperturbed HF wavefunction so that , and .
Using the equation of motion method [17], one obtains
where and
having introduced the symmetrized double commutator
The calculation of the block matrices yields
where ( and )
The occupation numbers are to be obtained from the one-body density matrix (OBDM)
The above system of nonlinear coupled equations is solved iteratively. Each iteration gives a new OBDM whose diagonalization yields a new set of occupation numbers () and defines a new natural orbital single-particle basis. The iteration ends when the convergence is reached.
The whole process implies a double iteration. One starts with solving the standard RPA eigenvalue problem, obtaining thereby the initial amplitudes and . These amplitudes together with the HF occupation numbers are used to determine the density matrices in Eqs. (17). The diagonalization of these latter quantities yields the new occupations numbers and a natural orbital basis to be used for solving again the RPA eigenvalue problem (Eq. (10)). A new set of RPA amplitudes is obtained thereby. The above procedure is iterated until the OBDM converges to a fixed point.
Final RPA amplitudes define excited states built on top of the correlated ground state and can be used to evaluate the ground-state transition amplitudes of any one-body operator
Note that in the limit the standard RPA expressions are recovered.
Numerical implementation and results. The Hamiltonian is composed of the intrinsic kinetic energy and the chiral potential N2LOGO(394) [30].
A HF basis was generated within an harmonic oscillator (HO) space covering all major shells up to . Such a basis was adopted to determine the matrix elements of the interaction. The NuHamil numerical code was used for this purpose [31].
All matrix elements of the three-body potential up to are included at the normal ordered two-body (NO2B) level [32]. These are the only ones entering in calculations performed within a p-h configuration space as in our case. Such a p-h basis, with the exception of the HF vacuum, is also free of spurious center of mass admixtures. These are removed by an orthogonalization procedure [33].
The ground-state energy is given by
where is the HF term and
comes from the correlations. The neutron () and proton () square radii
are referred to the center of mass (c.m.) in order to minimize the spurious admixtures present in the HF vacuum. The empirical charge radii are deduced from through the formula given, for instance, in Refs. [27, 34].
As shown in Fig. 1, the ground-state energies as well as the charge radii become insensitive to the HO frequency only in a space encompassing an increasing number of major shells as we move from light to heavy nuclei. An overall fair agreement between computed and empirical values is attained. For nuclei up to 90Zr the employed model space is large enough to yield stable results, being their dependence on the HO frequency very weak. Thus, the numerical results are presented for a single choice MeV.
The systematics presented in Fig. 2 prove that HF underestimates severely the binding energy of all studied nuclei. On the other hand, a strong over-binding is obtained once we add the contribution from the RPA ground-state correlations. The quenching action of the RRPA restores the consistency with the empirical values. The close agreement of the RRPA energies with the corresponding quantities obtained in HF + second-order perturbation theory (PT) is noteworthy.
RPA tends to overestimate also the nuclear radii (Fig. 2). RPA restores the agreement between theoretical and empirical quantities.
Next, we discuss the effect of renormalisation on selected excited states. In RPA, few levels of the 16O and 90Zr spectra are strongly pushed down toward the ground state (Fig. 3). This is a general feature occurring in all nuclei investigated. In some of them, like 40Ca, the lowest level collapses and the energy becomes imaginary. These instabilities are removed within the RRPA and the consistency with the experiments is restored.
Further insight can be obtained from the transition probabilities. The ground-state correlations, accounted for in RRPA, do not produce any significant further fragmentation of the strength with respect to RPA. On the other hand, because of the energy downshift and damping of the main E1 peak they cause, the energy-weighted sum rule, preserved in RPA, is sensibly underestimated (Fig. 4). This impact is to be attributed to the depletion of few single particle (hole) states since the majority of the states involved in the transitions, being far from the Fermi surface, are empty (full) (Fig.5).
The depletion of single-particle levels around such a surface affects strongly the low-lying octupole transitions (Table 1). In fact, the ground-state reduced strength of the transition to the 3, practically negligible in TDA, increases dramatically as we move to RPA which systematically overestimates the measured values. These are sensibly underestimated in RRPA as a consequence of the drastic quenching of the amplitudes (Fig. 5).
It is not difficult to find the reason of these dramatic changes in going from one approach to the other. Small variations in the and amplitudes are strongly amplified by their mutual interference produced once the transition amplitudes get squared.
| TDA | RPA | RRPA | Exp. | |
Concluding remarks. The limits of RPA emerge definitely from the present systematic based on the use of modern realistic potentials: The overestimation of binding energies and charge radii, the softness of some excited levels toward the ground state.
The origin of these deficiencies is suggested by the plot in Fig. 5. We infer from the too large values of the backward amplitudes that the strong ground-state correlations induced by modern realistic potentials invalidate the QBA underlying RPA. Once this approximation is removed, the amplitudes are damped and the inconsistencies removed.
The present approach achieves such a remarkable result without spoiling the simple structure of the RPA eigenvalue equations. One needs to account for the deviations of the single particle occupation numbers from 0 or 1. These deviations are quite pronounced in proximity of the Fermi surface and vanish rapidly as we move away from (Fig. 5). Moreover, it is necessary to include the contribution coming from the non-diagonal particle-particle (pp’) and hole-hole (hh’) (Eqs. 17) terms entering the OBDMs, as proposed in Ref. [22], in order to reproduce the charge radii, strongly overestimated in RPA.
As pointed out already, the RRPA levels and responses are referred to the true correlated ground state. A significant link with no-core shell model [39] and coupled-cluster [40] is established thereby. Such a link will be reinforced once the method will be recast in terms of Bogoliubov quasiparticles (qp), a feasible task, so as to cover a large fraction of open shell nuclei.
Even so reformulated, however, RRPA cannot be considered a complete alternative to no-core shell model. Being confined within a p-h or 2qp configuration space, RRPA can account only for a fraction of the energy levels forming the nuclear spectra and is unable to fully satisfy the energy-weighted sum rule. One needs to enlarge the space so as to include () or () configurations in order to cover the full energy spectrum and, hopefully, to reduce the persisting discrepancies between theoretical and experimental responses.
This upgrade is not at all trivial if we stick on RPA and its extensions. It is more natural to resort to a closely related and reliable alternative represented by the EMPM in its qp version [41].
We can therefore conclude that, within the limits of its validity, the present self-consistent RRPA approach offers, from first principles, a simple reliable unified systematic of bulk and spectroscopic properties of finite closed (sub-)shell nuclei, and is especially useful in the regions of medium and heavy nuclei, not easily accessible to more complete ab initio approaches.
Acknowledgements.
This work is supported by the Czech Science Foundation (Czech Republic), P203-23-06439S. Computational resources were provided by the CESNET LM2015042 and the CERIT Scientific Cloud LM2015085, under the program ”Projects of Large Research, Development, and Innovations Infrastructures”.References
- Rowe [2010] D. J. Rowe, Nuclear collective motion: models and theory (World Scientific, 2010).
- Ring and Schuck [2004] P. Ring and P. Schuck, The nuclear many-body problem (Springer Science & Business Media, 2004).
- Bender et al. [2003] M. Bender, P.-H. Heenen, and P.-G. Reinhard, Rev. Mod. Phys. 75, 121 (2003).
- Roca-Maza and Paar [2018] X. Roca-Maza and N. Paar, Progress in Particle and Nuclear Physics 101, 96 (2018).
- Vretenar et al. [2005] D. Vretenar, A. Afanasjev, G. Lalazissis, and P. Ring, Physics Reports 409, 101 (2005).
- Paar et al. [2003] N. Paar, P. Ring, T. Nikšić, and D. Vretenar, Phys. Rev. C 67, 034312 (2003).
- Lenske and Wambach [1990] H. Lenske and J. Wambach, Physics Letters B 249, 377 (1990).
- Litvinova [2023] E. Litvinova, The European Physical Journal A 59, 291 (2023).
- Gambacurta et al. [2015] D. Gambacurta, M. Grasso, and J. Engel, Phys. Rev. C 92, 034303 (2015).
- Paar et al. [2006] N. Paar, P. Papakonstantinou, H. Hergert, and R. Roth, Phys. Rev. C 74, 014318 (2006).
- Papakonstantinou and Roth [2010] P. Papakonstantinou and R. Roth, Phys. Rev. C 81, 024317 (2010).
- Hergert et al. [2011] H. Hergert, P. Papakonstantinou, and R. Roth, Phys. Rev. C 83, 064317 (2011).
- Wu et al. [2018] Q. Wu, B. S. Hu, F. R. Xu, Y. Z. Ma, S. J. Dai, Z. H. Sun, and G. R. Jansen, Phys. Rev. C 97, 054306 (2018).
- Raimondi and Barbieri [2019] F. Raimondi and C. Barbieri, Phys. Rev. C 99, 054327 (2019).
- Beaujeault-Taudière et al. [2023] Y. Beaujeault-Taudière, M. Frosini, J.-P. Ebran, T. Duguet, R. Roth, and V. Somà, Phys. Rev. C 107, L021302 (2023).
- Rowe [1968a] D. J. Rowe, Phys. Rev. 175, 1283 (1968a).
- Rowe [1968b] D. J. Rowe, Rev. Mod. Phys. 40, 153 (1968b).
- Ellis [1987] P. Ellis, Nuclear Physics A 467, 173 (1987).
- Karadjov et al. [1993] D. Karadjov, V. Voronov, and F. Catara, Physics Letters B 306, 197 (1993).
- Catara et al. [1994] F. Catara, N. Dinh Dang, and M. Sambataro, Nuclear Physics A 579, 1 (1994).
- Catara et al. [1996] F. Catara, G. Piccitto, M. Sambataro, and N. Van Giai, Phys. Rev. B 54, 17536 (1996).
- Catara et al. [1998] F. Catara, M. Grasso, G. Piccitto, and M. Sambataro, Phys. Rev. B 58, 16070 (1998).
- Roth and Paar [2007] P. P. R. Roth and N. Paar, Physical Review C 75, 014310 (2007).
- Schuck et al. [2021] P. Schuck, D. S. Delion, J. Dukelsky, M. Jemai, E. Litvinova, G. Röpke, and M. Tohyama, Physics Reports 929, 1 (2021).
- Minato [2016] F. Minato, Phys. Rev. C 93, 044319 (2016).
- Barton et al. [2021] M. Barton, P. Stevenson, and A. Rios, Phys. Rev. C 103, 064304 (2021).
- Hu et al. [2016] B. S. Hu, F. R. Xu, Z. H. Sun, J. P. Vary, and T. Li, Phys. Rev. C 94, 014303 (2016).
- Hergert et al. [2016] H. Hergert, S. Bogner, T. Moris, A. Schwenk, and K. Tsukiyama, Physics Reports 621, 165 (2016), ISSN 0370-1573.
- De Gregorio et al. [2017a] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017a).
- Jiang et al. [2020] W. G. Jiang, A. Ekström, C. Forssén, G. Hagen, G. R. Jansen, and T. Papenbrock, Phys. Rev. C 102, 054301 (2020).
- Miyagi [2023] T. Miyagi, The European Physical Journal A - Hadrons and Nuclei 59, 150 (2023).
- Miyagi et al. [2022] T. Miyagi, S. R. Stroberg, P. Navrátil, K. Hebeler, and J. D. Holt, Phys. Rev. C 105, 014302 (2022).
- De Gregorio et al. [2021] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Physics Letters B 821, 136636 (2021).
- De Gregorio et al. [2017b] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017b).
- [35] National nuclear data center database, URL https://www.nndc.bnl.gov/nudat3/.
- Angeli and Marinova [2013] I. Angeli and K. Marinova, Atomic Data and Nuclear Data Tables 99, 69 (2013), ISSN 0092-640X.
- Goriely et al. [2019] S. Goriely, P. Dimitriou, M. Wiedeking, T. Belgya, R. Firestone, J. Kopecky, M. Krtička, V. Plujko, R. Schwengner, S. Siem, et al., The European Physical Journal A 55 (2019), ISSN 1434-601X.
- Spear [1989] R. H. Spear, Atomic Data and Nuclear Data Tables 42, 55 (1989), ISSN 0092-640X.
- Barrett et al. [2013] B. R. Barrett, P. Navrátil, and J. P. Vary, Progress in Particle and Nuclear Physics 69, 131 (2013).
- Hagen et al. [2014] G. Hagen, T. Papenbrock, M. Hjorth-Jensen, and D. J. Dean, Rep. Prog. Phys. 77, 096302 (2014).
- De Gregorio et al. [2016] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Phys Rev. C 93, 044314 (2016).

### v2

@@
 It can be derived from the equation of motion method [39] or from the linear response theory [35].
-It was implemented within many different frameworks, some using non-relativistic (e.g.
+It is an essential tool for investigating nuclear spectra and response functions, including pygmy and giant resonances, as well as weak-interaction processes that are of significant importance in astrophysical contexts.
+RPA was implemented within many different frameworks, some using non-relativistic (e.g.
 Refs.
@@
 Most of the recipes for obviating this shortcoming were enumerated and developed long ago by Rowe [38, 37].
+A more updated list of papers devoted to this task is presented and discussed in a recent review [41].
 They consist in reintroducing the correlations into the ground state without spoiling the simple structure of the RPA eigenvalue equations.
@@
 [9, 8].
-It is worth mentioning a more recent approach using the realistic Argonne v18 potential [23], as well as a dressed (D) RPA study performed within a self-consistent Green’s function scheme [14].
-A variety of works aiming to go beyond RPA is presented and discussed in a recent review [24].
-Other approaches for computing ground-state correlations are available.
-We mention an extension of TDA [25], a time-dependent density matrix (TDDM) formalism [26], a many-body perturbative approach [27], in-medium similarity renormalization group (IMSRG) [28] and an equation of motion multiphonon method (EMPM) [29].
+The method so reformulated was applied to nuclei by Papakonstantiou et al.
+[32] using the UCOM potential derived from the Argonne V18 nucleon-nucleon (NN) interaction through a unitary transformation.
+The effect of the ground-state correlations on several multipole responses resulted to be modest.
+Several alternative approaches accounting for ground-state correlations are available.
+We mention an extension of TDA [26], a time-dependent density matrix (TDDM) formalism [3], a many-body perturbative approach [21], coupled-cluster(CC) [18], in-medium similarity renormalization group (IMSRG) [19], and an equation of motion multiphonon method (EMPM) [11].
 The need for restoring the ground-state correlations is also dictated by the fact that HF accounts only for a fraction of the binding energy of all closed-(sub)shell nuclei throughout the periodic table if modern realistic potentials are adopted.
@@
 [11].
-In the present work, we perform a systematic study of bulk properties, nuclear responses, and spectra of a large number of nuclei ranging from 4He to 208Pb.
+In the present work, we adopted the method developed in Ref.
+[8] and extended to nuclei in Ref.
+[32] to carry out a systematic study of bulk properties, nuclear responses, and spectra of a large number of closed (sub-)shell nuclei ranging from 4He to 208Pb.
 We use the chiral potential N2LOGO(394) [22], incorporating the three-body force, to generate a self-consistent HF basis and then solve the eigenvalue equations within TDA, RPA, and RRPA.
-To our knowledge, this is the first ab initio extension of RPA which removes the instabilities induced by the QBA.
-It accounts for the nuclear bulk properties and provides a unified scheme where energy levels and responses are referred to the true correlated (rather than unperturbed) ground state.
+To our knowledge, this is the first ab initio renormalized version of RPA that directly employs the state-of-art two- and three-body interactions without intermediate smoothing steps.
+As we shall show, it describes within a unified framework nuclear bulk properties with accuracy comparable to the one achieved by the CC and IMSRG methods and yields, within the limits imposed by 1p-1h space, complete spectra and responses referred to the correlated ground state.
+Moreover, it covers all sub-shell-closed nuclei across the nuclear chart, including the heavy systems.
+All these results were obtained at a computational cost significantly lower than the one spent by the available IMSRG solver [40].
+Thus, the present approach offers a new efficient manageable tool for a systematic study of low and high energy nuclear properties from first principles.
+It should have, therefore, an important impact on nuclear structure and nuclear astrophysics.
 The RRPA formalism. Following the procedure of Refs. [37, 9, 8] we consider states of the form
@@
 Final RPA amplitudes define excited states built on top of the correlated ground state and can be used to evaluate the ground-state transition amplitudes of any one-body operator
-Note that in the limit the standard RPA expressions are recovered.
+In the limit , the standard RPA formula is recovered.
 Numerical implementation and results. The Hamiltonian is composed of the intrinsic kinetic energy and the chiral potential N2LOGO(394) [22].
@@
 2 prove that HF underestimates severely the binding energy of all studied nuclei.
+Second-order perturbation theory (PT) refines the calculated energies, significantly reducing discrepancies with the experimental values.
 On the other hand, a strong over-binding is obtained once we add the contribution from the RPA ground-state correlations.
 The quenching action of the RRPA restores the consistency with the empirical values.
-The close agreement of the RRPA energies with the corresponding quantities obtained in HF + second-order perturbation theory (PT) is noteworthy.
-RPA tends to overestimate also the nuclear radii (Fig. 2). RPA restores the agreement between theoretical and empirical quantities.
-Next, we discuss the effect of renormalisation on selected excited states.
+A more detailed analysis (Table 1) shows that RRPA reproduces the binding energies with an overall accuracy comparable to the one achieved within the CC and IMSRG approaches.
+| RRPA | PT | IMSRG(2) | CC | Exp. | |
+| 124.7 | 124.4 | 126.2 | 127.5 | 127.62 | |
+| 168 | 164 | 167 | 169 | 168.96 | |
+| 338 | 338 | 339 | 346 | 342.05 | |
+| 416 | 410 | 415 | 420 | 416.00 | |
+| 634 | 624 | 629 | 639 | 641.55 | |
+| 773 | 764 | 768 | 782 | 783.90 | |
+| 807 | 793 | 804 | 818 | 825.30 | |
+| 1045 | 1021 | 1026 | 1043 | 1102.84 |
+| RRPA | CC | Exp. | |
+| 6.3 | 5.6 | 6.13 | |
+| 3.8 | 3.9 | 4.79 | |
+| 4.6 | 4.1 | 3.83 |
+RRPA restores also the agreement between theoretical and empirical radii overestimated in RPA (Fig. 2).
+Next, we discuss the effect of the renormalization on selected excited states.
 In RPA, few levels of the 16O and 90Zr spectra are strongly pushed down toward the ground state (Fig.
 3).
-This is a general feature occurring in all nuclei investigated.
+This is a general feature of all investigated nuclei.
 In some of them, like 40Ca, the lowest level collapses and the energy becomes imaginary.
 These instabilities are removed within the RRPA and the consistency with the experiments is restored.
+As shown in Table 2, the low lying levels, obtained using the same potential and model space, are reproduced with comparable accuracy in RRPA and CC [22].
+The RRPA monopole energy centroids are also comparable to the corresponding quantities obtained within recent LIT IMSRG/CC approaches [6].
 Further insight can be obtained from the transition probabilities.
-The ground-state correlations, accounted for in RRPA, do not produce any significant further fragmentation of the strength with respect to RPA.
+The ground-state correlations, accounted for in RRPA.
+do not further produce any significant fragmentation of the strength with respect to RPA.
 On the other hand, because of the energy downshift and damping of the main E1 peak they cause, the energy-weighted sum rule, preserved in RPA, is sensibly underestimated (Fig.
@@
 This impact is to be attributed to the depletion of few single particle (hole) states since the majority of the states involved in the transitions, being far from the Fermi surface, are empty (full) (Fig.5).
-The depletion of single-particle levels around such a surface affects strongly the low-lying octupole transitions (Table 1).
+The depletion of single-particle levels around such a surface strongly affects the low-lying octupole transitions (Table 3).
 In fact, the ground-state reduced strength of the transition to the 3, practically negligible in TDA, increases dramatically as we move to RPA which systematically overestimates the measured values.
@@
 This upgrade is not at all trivial if we stick on RPA and its extensions. It is more natural to resort to a closely related and reliable alternative represented by the EMPM in its qp version [14].
-We can therefore conclude that, within the limits of its validity, the present self-consistent RRPA approach offers, from first principles, a simple reliable unified systematic of bulk and spectroscopic properties of finite closed (sub-)shell nuclei, and is especially useful in the regions of medium and heavy nuclei, not easily accessible to more complete ab initio approaches.
+We can therefore conclude that, within the limits of its validity, the present self-consistent RRPA approach offers, from first principles, a simple reliable unified systematic of bulk and spectroscopic properties of finite closed (sub-)shell nuclei at a moderate numerical computational cost, and is especially useful in the regions of medium and heavy nuclei, not easily accessible to the other well established ab initio approaches.
+Hopefully, it might open a new route, based on the direct use of bar modern potentials, to the large community adopting mean field approaches.
 Acknowledgements.
-This work is supported by the Czech Science Foundation (Czech Republic), P203-23-06439S.
-Computational resources were provided by the CESNET LM2015042 and the CERIT Scientific Cloud LM2015085, under the program ”Projects of Large Research, Development, and Innovations Infrastructures”.References
-- Rowe [2010] D. J. Rowe, Nuclear collective motion: models and theory (World Scientific, 2010).
-- Ring and Schuck [2004] P. Ring and P. Schuck, The nuclear many-body problem (Springer Science & Business Media, 2004).
-- Bender et al. [2003] M. Bender, P.-H. Heenen, and P.-G. Reinhard, Rev. Mod. Phys. 75, 121 (2003).
-- Roca-Maza and Paar [2018] X. Roca-Maza and N. Paar, Progress in Particle and Nuclear Physics 101, 96 (2018).
-- Vretenar et al. [2005] D. Vretenar, A. Afanasjev, G. Lalazissis, and P. Ring, Physics Reports 409, 101 (2005).
-- Paar et al. [2003] N. Paar, P. Ring, T. Nikšić, and D. Vretenar, Phys. Rev. C 67, 034312 (2003).
-- Lenske and Wambach [1990] H. Lenske and J. Wambach, Physics Letters B 249, 377 (1990).
-- Litvinova [2023] E. Litvinova, The European Physical Journal A 59, 291 (2023).
-- Gambacurta et al. [2015] D. Gambacurta, M. Grasso, and J. Engel, Phys. Rev. C 92, 034303 (2015).
-- Paar et al. [2006] N. Paar, P. Papakonstantinou, H. Hergert, and R. Roth, Phys. Rev. C 74, 014318 (2006).
-- Papakonstantinou and Roth [2010] P. Papakonstantinou and R. Roth, Phys. Rev. C 81, 024317 (2010).
-- Hergert et al. [2011] H. Hergert, P. Papakonstantinou, and R. Roth, Phys. Rev. C 83, 064317 (2011).
-- Wu et al. [2018] Q. Wu, B. S. Hu, F. R. Xu, Y. Z. Ma, S. J. Dai, Z. H. Sun, and G. R. Jansen, Phys. Rev. C 97, 054306 (2018).
-- Raimondi and Barbieri [2019] F. Raimondi and C. Barbieri, Phys. Rev. C 99, 054327 (2019).
-- Beaujeault-Taudière et al. [2023] Y. Beaujeault-Taudière, M. Frosini, J.-P. Ebran, T. Duguet, R. Roth, and V. Somà, Phys. Rev. C 107, L021302 (2023).
-- Rowe [1968a] D. J. Rowe, Phys. Rev. 175, 1283 (1968a).
-- Rowe [1968b] D. J. Rowe, Rev. Mod. Phys. 40, 153 (1968b).
-- Ellis [1987] P. Ellis, Nuclear Physics A 467, 173 (1987).
-- Karadjov et al. [1993] D. Karadjov, V. Voronov, and F. Catara, Physics Letters B 306, 197 (1993).
-- Catara et al. [1994] F. Catara, N. Dinh Dang, and M. Sambataro, Nuclear Physics A 579, 1 (1994).
-- Catara et al. [1996] F. Catara, G. Piccitto, M. Sambataro, and N. Van Giai, Phys. Rev. B 54, 17536 (1996).
-- Catara et al. [1998] F. Catara, M. Grasso, G. Piccitto, and M. Sambataro, Phys. Rev. B 58, 16070 (1998).
-- Roth and Paar [2007] P. P. R. Roth and N. Paar, Physical Review C 75, 014310 (2007).
-- Schuck et al. [2021] P. Schuck, D. S. Delion, J. Dukelsky, M. Jemai, E. Litvinova, G. Röpke, and M. Tohyama, Physics Reports 929, 1 (2021).
-- Minato [2016] F. Minato, Phys. Rev. C 93, 044319 (2016).
-- Barton et al. [2021] M. Barton, P. Stevenson, and A. Rios, Phys. Rev. C 103, 064304 (2021).
-- Hu et al. [2016] B. S. Hu, F. R. Xu, Z. H. Sun, J. P. Vary, and T. Li, Phys. Rev. C 94, 014303 (2016).
-- Hergert et al. [2016] H. Hergert, S. Bogner, T. Moris, A. Schwenk, and K. Tsukiyama, Physics Reports 621, 165 (2016), ISSN 0370-1573.
-- De Gregorio et al. [2017a] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017a).
-- Jiang et al. [2020] W. G. Jiang, A. Ekström, C. Forssén, G. Hagen, G. R. Jansen, and T. Papenbrock, Phys. Rev. C 102, 054301 (2020).
-- Miyagi [2023] T. Miyagi, The European Physical Journal A - Hadrons and Nuclei 59, 150 (2023).
-- Miyagi et al. [2022] T. Miyagi, S. R. Stroberg, P. Navrátil, K. Hebeler, and J. D. Holt, Phys. Rev. C 105, 014302 (2022).
-- De Gregorio et al. [2021] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Physics Letters B 821, 136636 (2021).
-- De Gregorio et al. [2017b] G. De Gregorio, J. Herko, F. Knapp, N. Lo Iudice, and P. Veselý, Phys. Rev. C 95, 024306 (2017b).
-- [35] National nuclear data center database, URL https://www.nndc.bnl.gov/nudat3/.
-- Angeli and Marinova [2013] I. Angeli and K. Marinova, Atomic Data and Nuclear Data Tables 99, 69 (2013), ISSN 0092-640X.
-- Goriely et al.
-[2019] S.
-Goriely, P.
-Dimitriou, M.
-Wiedeking, T.
-Belgya, R.
-Firestone, J.
-Kopecky, M.
-Krtička, V.
-Plujko, R.
-Schwengner, S.
-Siem, et al., The European Physical Journal A 55 (2019), ISSN 1434-601X.
-- Spear [1989] R. H. Spear, Atomic Data and Nuclear Data Tables 42, 55 (1989), ISSN 0092-640X.
-- Barrett et al. [2013] B. R. Barrett, P. Navrátil, and J. P. Vary, Progress in Particle and Nuclear Physics 69, 131 (2013).
-- Hagen et al. [2014] G. Hagen, T. Papenbrock, M. Hjorth-Jensen, and D. J. Dean, Rep. Prog. Phys. 77, 096302 (2014).
-- De Gregorio et al. [2016] G. De Gregorio, F. Knapp, N. Lo Iudice, and P. Veselý, Phys Rev. C 93, 044314 (2016).
+This work is supported by the Czech Science Foundation (Czech Republic), P203-26-21972S.
+Computational resources were provided by the e-INFRA CZ project (ID:90254), supported by the Ministry of Education, Youth and Sports of the Czech Republic, and by the ELIXIR-CZ project (ID:90255).References
+- [1] (2013) Table of experimental nuclear ground state charge radii: an update.
+Atomic Data and Nuclear Data Tables 99 (1), pp.
+69 – 95.
+External Links: Document, ISSN 0092-640X Cited by: Figure 1, Figure 2.
+- [2] (2013) Ab initio no core shell model.
+Progress in Particle and Nuclear Physics 69, pp.
+131.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [3] (2021-06) Nuclear ground states in a consistent implementation of the time-dependent density matrix approach.
+Phys.
+Rev.
+C 103, pp.
+064304.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [4] (2023-02) Zero- and finite-temperature electromagnetic strength distributions in closed- and open-shell nuclei from first principles.
+Phys.
+Rev.
+C 107, pp.
+L021302.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [5] (2003-01) Self-consistent mean-field models for nuclear structure.
+Rev.
+Mod.
+Phys.
+75, pp.
+121–180.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [6] (2025) Ab initio calculations of monopole sum rules: from finite nuclei to infinite nuclear matter.
+arXiv:2511.15836.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [7] (1994) Ground-state correlations beyond rpa.
+Nuclear Physics A 579 (1), pp.
+1–12.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [8] (1998-12) Self-consistent determination of the one-body density matrix and particle-hole excitations.
+Phys.
+Rev.
+B 58, pp.
+16070–16075.
+External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [9] (1996) Towards a self-consistent random-phase approximation for fermi systems.
+Phys.
+Rev.
+B 54 (24), pp.
+17536–17546.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [10] Data extracted using the NNDC On-line Data Service from the ENSDF database. Cited by: Table 1.
+- [11] (2017-02) Ground-state correlations within a nonperturbative approach.
+Phys.
+Rev.
+C 95, pp.
+024306.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [12] (2017-02) Ground-state correlations within a nonperturbative approach.
+Phys.
+Rev.
+C 95, pp.
+024306.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [13] (2021) Removal of the center of mass in nuclei and its effects on 4he.
+Physics Letters B 821, pp.
+136636.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [14] (2016) Self-consistent quasiparticle formulation of a multiphonon method and its application to the neutron-rich nucleus.
+Phys Rev.
+C 93, pp.
+044314.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [15] (1987) Particle-hole and particle-particle rpa ground state correlations.
+Nuclear Physics A 467 (2), pp.
+173–184.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [16] (2015-09) Subtraction method in the second random-phase approximation: first applications with a skyrme energy functional.
+Phys.
+Rev.
+C 92, pp.
+034303.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [17] (2019) Reference database for photon strength functions. The European Physical Journal A 55 (10). External Links: ISSN 1434-601X, Document Cited by: Figure 4.
+- [18] (2014) Coupled-cluster computations of atomic nuclei.
+Rep.
+Prog.
+Phys.
+77, pp.
+096302.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [19] (2016) The in-medium similarity renormalization group: a novel ab initio method for nuclei.
+Physics Reports 621, pp.
+165–222.
+External Links: ISSN 0370-1573 Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [20] (2011-06) Quasiparticle random-phase approximation with interactions from the similarity renormalization group.
+Phys.
+Rev.
+C 83, pp.
+064317.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [21] (2016) Ab initio nuclear many-body perturbation calculations in the hartree-fock basis.
+Phys.
+Rev.
+C 94, pp.
+014303.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [22] (2020) Accurate bulk properties of nuclei from to from potentials with isobars.
+Phys.
+Rev.
+C 102, pp.
+054301.
+Cited by: Table 1, Table 2, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [23] (1993) Ground state correlations and charge transition densities.
+Physics Letters B 306 (3), pp.
+197–200.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [24] (1990) RPA ground state correlations in nuclei.
+Physics Letters B 249 (3), pp.
+377–380.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [25] (2023) On the dynamical kernels of fermionic equations of motion in strongly-correlated media.
+The European Physical Journal A 59 (12), pp.
+291.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [26] (2016-04) Estimation of a 2p2h effect on gamow-teller transitions within the second tamm-dancoff approximation.
+Phys.
+Rev.
+C 93, pp.
+044319.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [27] (2022-01) Converged ab initio calculations of heavy nuclei.
+Phys.
+Rev.
+C 105, pp.
+014302.
+External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [28] (2023) NuHamil. Eur. Phys. J. A 59, pp. 150. External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [29] National nuclear data center database. External Links: Link Cited by: Figure 1, Figure 2.
+- [30] (2006-07) Collective multipole excitations based on correlated realistic nucleon-nucleon interactions.
+Phys.
+Rev.
+C 74, pp.
+014318.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [31] (2003-03) Quasiparticle random phase approximation based on the relativistic hartree-bogoliubov model.
+Phys.
+Rev.
+C 67, pp.
+034312.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [32] (2007-01) Nuclear collective excitations using correlated realistic interactions: the role of explicit random-phase approximation correlations.
+Phys.
+Rev.
+C 75, pp.
+014310.
+External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [33] (2010-02) Large-scale second random-phase approximation calculations with finite-range interactions.
+Phys.
+Rev.
+C 81, pp.
+024317.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [34] (2019-05) Nuclear electromagnetic dipole response with the self-consistent green’s function formalism.
+Phys.
+Rev.
+C 99, pp.
+054327.
+External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [35] (2004) The nuclear many-body problem.
+Springer Science & Business Media.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [36] (2018) Nuclear equation of state from ground and collective excited state properties of nuclei.
+Progress in Particle and Nuclear Physics 101, pp.
+96.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [37] (1968-01) Equations-of-motion method and the extended shell model.
+Rev.
+Mod.
+Phys.
+40, pp.
+153–166.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [38] (1968-11) Methods for calculating ground-state correlations of vibrational nuclei.
+Phys.
+Rev.
+175, pp.
+1283–1292.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [39] (2010) Nuclear collective motion: models and theory. World Scientific. Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [40] GitHub. Note: https://github.com/ragnarstroberg/imsrg Cited by: Table 1, Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [41] (2021) Equation of motion method for strongly correlated fermi systems and extended rpa approaches.
+Physics Reports 929, pp.
+1–84.
+External Links: Document Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [42] (1989) Reduced electric-octupole transition probabilities, B(E3; ), for even-even nuclides throughout the periodic table.
+Atomic Data and Nuclear Data Tables 42 (1), pp.
+55–104.
+External Links: ISSN 0092-640X, Document Cited by: Table 3.
+- [43] (2005) Relativistic hartree–bogoliubov theory: static and dynamic aspects of exotic nuclear structure.
+Physics Reports 409 (3), pp.
+101.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.
+- [44] (2018-05) Chiral descriptions of nuclear multipole resonances within the random-phase approximation.
+Phys.
+Rev.
+C 97, pp.
+054306.
+Cited by: Bulk and spectroscopic nuclear properties within an ab initio renormalized random-phase approximation framework.